OpenStudy (anonymous):

pls help this is my last question!!! definite integral from -1/2 to 1/2 x^2/square root of 1-x

OpenStudy (anonymous):

\[\huge \int\limits_{-\frac{1}{2}}^\frac{1}{2}\frac{x^2}{\sqrt{1-x}}dx\] Try substituting \(u=1-x\).

OpenStudy (anonymous):

if u=1-x , du=-1 then nothing

OpenStudy (anonymous):

Did you substitute into the integrand?

OpenStudy (anonymous):

yes it does not work

OpenStudy (anonymous):

Yes it does...distribute the numerator.

OpenStudy (anonymous):

:( i do not know how to do it can you explain

OpenStudy (anonymous):

\[\frac{(1-u)^2}{\sqrt u}=\frac{1-2u+u^2}{\sqrt u}=u^{-\frac{1}{2}}-2u^{\frac{1}{2}}+u^{\frac{3}{2}}\]